A poser has a length 4 inches greather than is width w. Demetrius will place a frame 2 inches wide around the poster. The area of the poster and frame together is 672 square inches. Which factored equation could be used to find the width of the poster?
Answer choices:
A) (w+32(w-20)=0
B) (w+4)(w+8)=0
C) (w+2)(w+6)=672
D) (w-32)(w+20)=0

Answer by uma(370)

Let the width of the poster be w
Then length = (w+4)
When there is a frame of 2 inches wide,
length of the frame = (w+4) +2 + 2
= w + 8

Width of the frame = (w) + 2+ 2
= w + 4
Area = length * width
==> (w+8)(w+4) = 672

Good Luck!!!